### v1

# LES-Talker: Fine-Grained Emotion Editing for Talking Head Generation in Linear Emotion Space

Abstract
While existing one-shot talking head generation models have achieved progress in coarse-grained emotion editing, there is still a lack of fine-grained emotion editing models with high interpretability. We argue that for an approach to be considered fine-grained, it needs to provide clear definitions and sufficiently detailed differentiation. We present LES-Talker, a novel one-shot talking head generation model with high interpretability, to achieve fine-grained emotion editing across emotion types, emotion levels, and facial units. We propose a Linear Emotion Space (LES) definition based on Facial Action Units to characterize emotion transformations as vector transformations. We design the Cross-Dimension Attention Net (CDAN) to deeply mine the correlation between LES representation and 3D model representation. Through mining multiple relationships across different feature and structure dimensions, we enable LES representation to guide the controllable deformation of 3D model. In order to adapt the multimodal data with deviations to the LES and enhance visual quality, we utilize specialized network design and training strategies. Experiments show that our method provides high visual quality along with multilevel and interpretable fine-grained emotion editing, outperforming mainstream methods. Project page: https://peterfanfan.github.io/LES-Talker/
Figure 1: Linear Emotion Space (LES) based on Facial Action Units (AUs) supports our LES-Talker model, offering exceptional interpretability. It enables fine-grained editing across 8 emotion types, 17 facial units, and continuous levels above 0. It can be driven by a lightweight use of video (requiring a sequence of images to provide the AU source) or by audio alone.
11footnotetext: These authors contributed equally.22footnotetext: * Corresponding author.
1 Introduction
Talking head generation has attracted significant attention from researchers in recent years and has many applications in the fields of digital human creation [45], virtual reality [7], etc. To enhance the diversity and expressiveness of talking head generation, emotion editing tasks were emphasized. However, existing studies have the following two notable drawbacks: (1) the interpretability of methods is lacking; (2) the effectiveness of fine-grained emotion editing is still unsatisfactory. On the one hand, some studies offer expression reference images without clear emotion definition [33, 34, 20], while others employ discrete emotion labels [17, 32, 24]. Additionally, some extract emotion features from latent spaces [34, 25] leading to an implicit transformation of emotion. Although Facial Action Units (AUs) effectively describe emotions, various methods [13, 16, 14] use different combinations of AUs for the same emotion. Moreover, AUs alone remain insufficient. On the other hand, several studies [35, 25, 33] focus on generating videos with specific emotions, realizing only coarse-grained emotion editing. One of the few fine-grained emotion studies [34, 31] also shows only discrete emotion levels in three emotion types.
We argue that for an approach to be considered fine-grained, it needs to offer clear definitions and detailed differentiation. Motivated by the above, we propose Linear Emotion Space (LES), a definition that explicitly characterizes emotion transformations. LES supports us in proposing LES-Talker, a video generation model for editing multiple emotion types, continuous emotion levels, and individual Facial Action Units. To provide an interpretable theoretical foundation, we develop an LES definition based on elements derived from AUs. Using a coarse-to-fine emotion strategy, LES characterizes emotion transformations as vector changes in two subspaces: the Action Subspace, representing facial unit actions, and the Isolation Subspace, capturing subtle emotion details. When fine-grained emotion editing is required, the vectors in the LES can be used as intermediate representations of emotions. Specifically, when editing a neutral emotion video, each frame is extracted with AUs values and mapped into the Action and Isolation Subspaces as vector representations. The target emotion with specified level corresponds to a vector in these spaces, allowing fine-grained emotion transformations through vector transformations. Further, each dimension of LES has a physical counterpart, making the meaning of these edits explicit.
To achieve a satisfactory editing effect, our LES-Talker utilizes LES vectors and 3DMM coefficients as two intermediate representations. We divide the task into three major components. First, we aim to adapt the multi-modal data to the LES definition. We transform the AUs sequences, enhance audios and optimize predicted 3D coefficients, through direct adaptation, wav2vec-based [2] audio encoder, and residual learning, respectively. After generating the accurate representations in LES, we apply vector operation to inject emotion according to the definition in LES. We also propose a Cross-Dimension Attention Net (CDAN). By combining CDANs in series and parallel, we enable detailed guidance of the LES representation in the deformation of the 3D model. Inspired by Sadtalker [42], we use the identity and pose information from reference images and edited 3D coefficients to generate a video through a novel render.
Our study makes the following four main contributions:
  • We propose the Linear Emotion Space (LES), a fine-grained emotion definition that provides an interpretable theoretical foundation for fine-grained emotion research.
  • We introduce LES-Talker, a novel one-shot talking head generation model with high interpretability, designed to achieve fine-grained emotion editing across emotion types, levels, and facial units.
  • We propose a novel and universal Cross-Dimension Attention Network to mine potential correlations between 3D model representation and LES representation, which enables the detailed guidance of LES representation in the deformation of the 3D model.
  • Quantitative and qualitative experiments demonstrate our method generates videos with high visual quality and is capable of fine-grained editing at multiple levels.
Figure 2: Pipeline of LES-Talker. Inputs include an identity image, audio, optional AU Source, and user editing targets. The Linear Emotion Space Recon. generates emotion vectors in LES. Emo Injector transforms these vectors based on user targets $(emo,level)\text{ or }(au,bias)$. Two levels of Cross-Dimension Attention Net (CDAN) process decomposed vectors $\boldsymbol{u}$ and $\boldsymbol{v}$ to produce 3D coefficients, optimized via Offset Decoder. These coefficients, along with identity information, create the rendered video.
2 Related work
2.1 Talking Head Video Generation
Existing research on talking head generation is mainly categorized into audio-driven [36, 41, 30, 1, 19, 6] and video-driven [23, 37, 12, 46] two broad categories. In audio-driven methods, lightweight use of video to compensate the lack of information is a common approach, such as StyleTalk [24] extracting facial motion patterns using a transformer, and PC-AVS [48] incorporating head pose information through implicit low-dimensional pose coding. Video-driven methods utilize the richer information contained in the input video to generate more natural results, focusing on different driving methods and intermediate representations, which are divided into implicit and explicit. Implicit modeling represents scenes by learning mathematical functions, such as Signed Distance Functions (SDF) [38, 47, 29] and Neural Radiation Fields (NeRF) [18, 21, 28]. Explicit modeling directly constructs editable 3D geometric representations, such as 3D Morphable Model (3DMM) [42, 27, 39] and 3D Gaussian Splatting (3D-GS) [40, 9, 22].
We adopt LES vectors and 3DMM coefficients as representations for explicit control methods.
2.2 Fine-grained Facial Emotion Video Generation
In talking head generation, early studies [26, 8] focused on improving visual quality, while recent efforts have explored coarse-grained facial emotion editing. EmoTalk [25] employs an audio-driven method using two different audio extractors to separate emotion and content, and EDTalk [33] emphasizes controlled talking head generation by decoupling mouth shape, head pose, and emotional expression. However, these studies only achieve coarse-grained emotion expression.Recently, a few fine-grained emotion studies have emerged. Under co-driven conditions, PD-FGC [34] demonstrates linear interpolation between expression features from different sources by learning the expression latent space. EmoSpeaker [15] decouples audio input into content and emotion vectors, editing content vectors for different emotions. FG-EmoTalk [31] shows video-driven realization of three defined emotions and several levels of fine-grained emotions using AUs as control inputs. However, these studies lack clarity on the principles of fine-grained emotion changes and primarily realize discrete rather than continuous changes.
While utilizing AUs for fine-grained emotions is promising, it is insufficient. Our study refines elements derived from AUs, providing a rigorous definition and demonstrating talking head generation with continuously varying level and multiple emotion options, allowing individual control of specific facial units.
3 Method
We present the Linear Emotion Space (LES) and LES-Talker that aim to achieve fine-grained emotion editing with exceptional interpretability. The complete pipeline of our one-shot method is depicted in Fig. 2. Below, we first provide a brief introduction to Action Units (AUs) and 3D face models as preliminaries in Sec. 3.1. Then, in Sec. 3.2, we offer a precise and rigorous definition of the Linear Emotion Space, which serves as the theoretical foundation of our model. Cross-Dimension Attention Network (CDAN) plays a key role in enabling the LES representation to guide the controllable deformation of the 3D model, and we detail its design in Sec. 3.3. Finally, in Sec. 3.4, we describe the design of other key components that adapt multi-modal features to the LES and improve visual quality.
3.1 Preliminary
Facial Action Units. Facial Action Units (AUs) are essential in studying facial expressions. AUs describe patterns of facial muscle movements associated with emotions. Currently, there are seventeen controllable AUs. For instance, in happiness, relevant AUs include AU6 (cheek muscle contraction, forming a smile), AU12 (lip corner elevation, indicating pleasure), AU25 (mouth opening, signifying excitement), and AU4 (brow raising, expressing joy or surprise). Different AUs can be used individually or in combination to represent emotional states; however, varying definitions across studies can cause confusion. Our method utilizes the complete set of AUs, along with subtle information beyond AUs, to represent each emotional state. Further explanations of AUs are in the supplementary materials.
3D Morphable Model. We use 3D Morphable Models (3DMMs) as one of our intermediate representations, inspired by the single image deep 3D reconstruction method [11] and recent talking head generation method [42]. The 3D face shape S and the talking head motion M can be decoupled as:
$\begin{array}[]{c}S=\overline{S}+\alpha\cdot U_{\text{id}}+\beta\cdot U_{\text{exp}}\\ M=[\beta,\gamma]\end{array}$ (1)
$\overline{S}$ is the average shape of the 3D face. $U_{\text{id}}$ and $U_{\text{exp}}$ are the orthonormal basis [4]. Coefficients $\alpha\in\mathbb{R}^{80}$, $\beta\in\mathbb{R}^{64}$ and $\gamma\in\mathbb{R}^{6}$ describe the person identity, face expression and head pose respectively.
3.2 Linear Emotion Space
As mentioned in the Preliminary section, there are 17 controllable AUs, indexed from 1 to 17. Their values could be extracted from the emotional frame. Through two different optimization functions ($\text{Opt}_{\text{1}}$ and $\text{Opt}_{\text{2}}$), we derive a point $\boldsymbol{w}$ with 41 coordinates, each having a specific physical meaning. In the following descriptions, points, and vectors in the LES are equivalent.
The Linear Emotion Space is based on two hypotheses: (1) The space containing $\boldsymbol{w}$ is linear. (2) The level of emotion $int_{emo}$ varies linearly in this space. Therefore, the Linear Emotion Space $\mathbb{E}$ is defined with the following characteristics:
$if\ \boldsymbol{w}=(e_{1},e_{2},\ldots,e_{41})\in\mathbb{E}\quad then\ \boldsymbol{w}\in\mathbb{R}^{41}$ (2)
$e_{i}=\text{Opt}_{\text{1}}(AU_{i})\quad i\in[1,17]$ (3)
$e_{i}=\text{Opt}_{\text{2}}(AU_{j})\quad j=i-17\quad i\in[18,34]$ (4)
$e_{i}=0\text{ except }e_{index}=od\quad i\in[35,41]$ (5)
$\mathbb{A}=\{\boldsymbol{u}=(e_{1},\ldots,e_{17})\}\subseteq\mathbb{E}$ (6)
$\mathbb{I}=\{\boldsymbol{v}=(e_{18},\ldots,e_{41})\}\subseteq\mathbb{E}$ (7)
$level_{emo}\ \text{varies linearly in}\ \mathbb{E}$ (8)
Figure 3: Subspaces of Linear Emotion Space
Action Subspace. As shown in Fig. 3, the space $\mathbb{E}$ comprises two subspaces: the Action Subspace $\mathbb{A}$ and the Isolation Subspace $\mathbb{I}$.
Subspace $\mathbb{A}$ provides a representation for a unified AU level and emotion level by defining feature vectors for different emotional levels. The vector $\boldsymbol{u}\in\mathbb{A}$, as defined in Eqs. 3 and 6, primarily describes the actions of Facial Units. Considering the imbalanced distribution between AUs, we apply optimization function $\text{Opt}_{\text{1}}$:
$\text{Opt}_{\text{1}}(AU_{i})=\frac{AU_{i}-\mu_{D}}{\sigma_{D}}$ (9)
where $\mu_{D}$ and $\sigma_{D}$ are the mean and standard deviation of $AU_{i}$ based on the entire dataset. For each emotion, the Action Subspace defines three special $\boldsymbol{u}$ as feature vectors $\boldsymbol{uf}_{emo,level}$. The MEAD dataset classifies emotions excluding neutral into three base levels: $\text{level}_{\text{1}}\text{, level}_{\text{2}}\text{, and }\text{level}_{\text{3}}$, used as base level anchors. We individually collect $\boldsymbol{u}$ for each emotion at different levels to obtain $\boldsymbol{uf}_{emo,level}$:
$\boldsymbol{uf}_{emo,k}=\frac{1}{N}\sum\boldsymbol{u}_{emo,level_{k}}$ (10)
$\boldsymbol{uf}_{neutral,0}=\frac{1}{N}\sum\boldsymbol{u}_{neutral,level_{1}}$ (11)
Figure 4: Part of the Outlier Matrix.
Effectiveness Proof For a $\boldsymbol{uf_{emo,level}}$ to effectively capture the distinction from the general situation, it should exhibit outlier values. In the same video, the correlation between $\boldsymbol{u}$ decreases rapidly as the frame interval increases, whereas in different videos, $\boldsymbol{u}$ is independent. Belonging to the same AU, the $e_{i}$ is identically distributed. We apply the Central Limit Theorem and hypothesis testing to determine if an $e_{i}$ in $\boldsymbol{uf_{emo,int}}$ is an outlier. Given the parameters: $\mu=0$, $\sigma=1$, $n=45,000$, confidence level = 0.999 (Z = 3.291). An $e_{i}$ is a significant anomaly if:
$|e_{i}|>\frac{Z\cdot\sigma}{\sqrt{n}}=0.0155$ (12)
In the outlier matrix Fig. 4, most AUs corresponding to $e_{i}$ are outliers. The effectiveness has been proven.
Isolation Subspace. Subspace $\mathbb{I}$ provides a representation for subtle emotion transformations by isolating each emotional vector to capture information beyond AUs.
The vector $\boldsymbol{v}\in\mathbb{I}$ is defined in Eq. 7. As defined in Eq. 5, elements $e_{35}$ to $e_{41}$ use a one-hot-like encoding to represent seven emotions excluding neutral, and $index$ corresponds to an emotion type. The origin distance is calculated as:
$od=\sqrt{\sum_{i=18}^{34}{e_{i}^{2}}}$ (13)
This distance $od$ reflects emotion level tendency, which we leverage to enable the network (CDAN) to learn additional representations. To unify the tendency expressed by AUs, we apply the optimization function $\text{Opt}_{\text{2}}$:
$\text{Opt}_{\text{2}}(AU_{i})=\frac{|AU_{i}|}{\sigma_{emo}}$ (14)
where $\sigma_{emo}$ is the standard deviation of $AU_{i}$ for a specific emotion.
Isolation Proof For a vector $\boldsymbol{v}$ to effectively represent the emotion level tendency, it must be distinguishable from vectors of other emotions. Poor results from Gaussian and k-means clustering (Silhouette Scores: -0.0217 and 0.0137) in $\mathbb{A}$ indicate the necessity of emotion isolation within $\mathbb{I}$. Define both outer (Opd) and inner (Ipd) emotion distances:
$d=\|\boldsymbol{v}_{1}-\boldsymbol{v}_{2}\|$ (15)
For $\boldsymbol{v}$ representing different emotions but with the same Od:
$Ipd\leq od\times\sqrt{2}\leq Opd$ (16)
This condition ensures that the network is always able to isolate vectors of different emotions as the level increases. The isolation has been proven.
Figure 5: Structure of the Cross-Dimension Attention Net. Illustration in the process for a single frame’s coefficients. G denotes the vector outer product operation, and C denotes vector concatenation. $\boldsymbol{u}$,$\boldsymbol{v}$ are vectors in ACT and ISO Subspace respectively. $\beta$ is 3DMM coefficients.
3.3 Cross-Dimension Attention Net
Cross-Dimension Attention Net (CDAN) mines high-dimensional correlations between two low-dimensional inputs using matrix-attention and cross-dimension mechanism. The network inputs are two vectors, $\beta$ and either $\boldsymbol{u}$ or $\boldsymbol{v}$. Specifically, $\boldsymbol{u}\in\mathbb{A}$ is a $1\times 17$ vector, and $\boldsymbol{v}\in\mathbb{I}$ is a $1\times 24$ vector. The vector $\beta$, from Eq. 1, is a $1\times 64$ vector.
From a cross-dimensional perspective, Action Units (AUs) affect 3DMM locations differently; for example, chin movements impact the lips more than the eyes. Different facial regions correlate to varying degrees. CDAN explores these correlations in higher and lower dimensions.
For higher dimensions, vectors $\boldsymbol{u}$ and $\boldsymbol{\beta}$ are combined into the Initial Joint Coefficient Matrix $\boldsymbol{J}$:
$\boldsymbol{J}=\text{Combine}(\boldsymbol{u},\boldsymbol{\beta})$ (17)
Passing $\boldsymbol{J}$ through a linear layer produces the Combined Joint Coefficient Matrix $\boldsymbol{J^{*}}$:
$\boldsymbol{J^{*}}=\text{Linear}(\boldsymbol{J})$ (18)
Here, $\boldsymbol{J^{*}}$ is shaped $17\times 64$. The 17 dimensions represent facial units as sequence length, and the 64 dimensions represent 3DMM coefficients as embedding vectors. A matrix-attention mechanism adds to correlation mining.
For lower dimensions, concatenate $\boldsymbol{u}$ and $\beta$:
$\text{Concat}(\boldsymbol{u},\boldsymbol{\beta})\rightarrow\text{FC}(\cdot)$ (19)
Eq. 23 describes parallel inputs to predict transformed 3DMM coefficients:
$\beta^{\prime}=\text{MLP}(\text{FC}(\boldsymbol{u},\beta),\text{Att}(\boldsymbol{u},\beta))$ (20)
As shown in Fig. 2, utilizing $\boldsymbol{{w}^{{}^{\prime}}}\in\mathbb{E}$ from the Emo Injector, our method employs two CDANs in series and parallel to achieve controllable deformation of a facial 3D model. Each CDAN receives vectors from the ACT and ISO subspaces, respectively. For finer representation in the ISO subspace, when $\boldsymbol{v}$ is an input, the other input is the predicted coefficients $\beta^{\prime}$ obtained from $\boldsymbol{u}$ and $\beta$.
Figure 6: The Emotion Injector.
Figure 7: Fine-grained emotion editing.
3.4 Other Key Components
Linear Emotion Space Recon. The reconstruction of emotion utilizes OpenFace [3] to extract the origin AUs sequence from a given AU source and then processes the AUs sequence, mapping it to the ACT and ISO subspaces. This involves deriving $\mu_{D}$, $\sigma_{D}$, and each $\sigma_{emo}$ statistically before training, as shown in Eq. 9 and Eq. 14.
Audio Encoder-Decoder. Considering that it is not always easy for users to provide AU source video in actual use, our method can be video-driven as well as audio-driven. Shown in Fig. 2, an Audio Encoder, based on a ResNet model [42], encodes audio inputs into content vectors for further collaboration with Decoders. AU Decoder provides generative input for vector $\boldsymbol{w}\in\mathbb{E}$ when AU inputs are absent. Offset Decoder enhances the predicted 3D coefficients by incorporating content vectors and leveraging predictions from another pretrained audio encoder [42]. Since audio features contain unique information, Offset Decoder compensates for residuals in the 3D coefficients after CDAN prediction.
Emotion Injector. The Emotion Injector, based on Fig. 4, Eq. 9, and Eq. 10, utilizes twenty-two feature vectors $\boldsymbol{uf}$ for eight emotions, stored statistically before testing. $\boldsymbol{uf_{neutral,0}}$ is defined where emotion level is 0. $\boldsymbol{u}$ is updated first and then $\boldsymbol{v}$ is updated. As shown in Fig. 6 and based on Eq. 8, when user’s control label is $(emo,level)$, the transformation is as follows:
$\begin{array}[]{c}i=\lceil level\rceil\quad\quad\quad j=\lfloor level\rfloor\end{array}$ (21)
$\boldsymbol{uf}_{emo,level}=$ $(\boldsymbol{uf}_{emo,i}-\boldsymbol{uf}_{emo,j})$ (22)
$\cdot(level-j)+\boldsymbol{uf}_{emo,j}$
$\begin{array}[]{c}\boldsymbol{u}_{inj}=\boldsymbol{uf}_{emo,level}-\boldsymbol{uf}_{neutral,0}\end{array}$ (23)
$\begin{array}[]{c}\boldsymbol{u^{\prime}}=\boldsymbol{u}_{inj}+\boldsymbol{u}\end{array}$ (24)
3D-Aware Face Render. Inspired by Sadtalker [42], we use identity, pose, and transformed 3D coefficients to generate a video with a pre-trained renderer.
4 Experiment
In this section, we demonstrate three key points through multiple experiment settings: (1) the adaptation to Linear Emotion Space (LES) definition enables LES-Talker to achieve fine-grained emotion editing; (2) the improved visual achieved by LES-Talker; (3) the effective design of the CDAN and other key components.
Figure 8: Fine-grained AUs editing.
4.1 Experimental Settings.
Datasets. We utilized the MEAD [35] as our primary dataset. To demonstrate the generalizability of our method, we also conducted evaluations using the CREMA-D [5], HDTF [44], and a diverse dataset [43].
Implementation Details. We performed all experiments on an NVIDIA RTX 3090Ti GPU using the PyTorch platform. The input audio was sampled at 16,000 Hz and represented as a 0.2-second mel spectrogram. Videos were cropped and resized to 512$\times$512. OpenFace [3] and DeepFace3DReconstruction [11] were used to extract AU values and 3DMM coefficients from input images. Our training employs a two-step coarse-to-fine strategy. We provide more details about training in the supplementary materials.
Evaluation Metrics. Video quality was assessed using the following four metrics: Frechet Inception Distance (FID), Structural Similarity (SSIM), Peak Signal-to-Noise Ratio (PSNR), Cumulative Probability Blur Detection (CPBD). Lip synchronization was evaluated with three metrics using Syncnet [10]: Lip Sync Confidence (AVConf), Lip Offset (AVOffset), and Minimum Offset (MinDist).
Figure 9: t-SNE Visualization.
4.2 Editing Effectiveness
We demonstrate the effectiveness of the editing by achieving the expected experimental results.
Fine-grained emotion editing. In the test set, we selected identity images and AU sources from neutral emotion videos, along with the corresponding audio. The Emo Injector method was strictly followed to inject different emotions at various levels. We compared frames with the same emotion but varying levels. In Fig. 7, the horizontal axis represents the level, while the vertical axis represents different emotions. We achieved fine-grained editing for levels 0-3 and effective editing for hyperdomain levels (level $>$ 3).
Fine-grained AUs editing As shown in Fig. 1, we can set the level in a arbitrary values. To validate the model’s generalization and effectiveness, we imposed strict conditions: absent AU source, cross-dataset, and cross-identity. Control conditions involved injecting +2.5 and -2.5 into the AUs in the ACT Subspace, with each image pair having +2.5 on the left and -2.5 on the right. As illustrated in Fig. 8, our method achieves editing control over nearly all defined AUs.
Space Visualization. We generated 1,755 videos across five emotions, with levels from 0 to 4 in steps of 0.0114. Each video’s frames were transformed into 41-dimensional LES representations and averaged, followed by t-SNE for dimensionality reduction. In Fig. 9, we can observe that the spatial representation reveals a clear emotional gradient and distinct separation between emotions.
Figure 10: Visual quality comparison.
Method Video Quality Comparison Lip Synchronization Comparison User Study
FID↓ SSIM↑ PSNR↑ CPBD↑ MinDist↓ AVConf↑ AVOffset(→0) LipSyn↑ EmoAcc↑ Reality↑ Quality↑
Real Video 0.000 1.000 31.734 0.265 7.869 6.564 -2.000 4.54 4.74 4.88 4.89
MEAD 146.454 0.469 14.864 0.191 11.957 2.674 -2.000 2.73 3.42 3.87 3.72
EVP 56.650 0.453 16.308 0.341 12.443 3.163 5.000 2.83 3.55 3.71 3.85
EAMM 204.002 0.396 12.832 0.135 10.091 3.046 -4.000 3.12 3.72 3.65 3.56
SadTalker 25.746 0.718 20.728 0.249 8.875 6.484 1.000 3.71 - 3.30 3.86
Ours 24.783 0.743 21.686 0.261 9.800 7.284 0.000 3.59 3.79 3.96 4.10
Table 1: Quantitative comparisons on video quality, lip synchronization, and user study.
4.3 Visual Quaility
Comparision with Other Methods. To the best of our knowledge, no other research in talking head generation explicitly defines fine-grained emotions. We focus on visual quality, which is tested randomly in the neutral generation case of the MEAD dataset. Merely generating lip movements is insufficient for real-world applications. We compare our method with state-of-the-art emotion-driven facial expression generation methods, such as EAMM, EVP, and MEAD, as well as the one-shot generation method SadTalker. Our method demonstrates improved visual quality in Fig. 10. Tab. 1 shows our method exhibits superior performance in video quality and lip synchronization.
Emo Type Emo Level
0.50 1.00 1.50 2.00 2.50
Angry 0.87 1.20 1.60 1.80 1.84
Happy 0.83 1.41 1.77 2.00 2.23
Sad 1.22 1.42 1.51 2.07 2.07
Surprised 1.73 1.92 2.01 2.04 2.28
Average 1.16 1.49 1.72 1.98 2.10
Table 2: User-perceived level.
User Study. We conducted user studies with 15 participants to evaluate our model across multiple metrics (rated 1–5) compared to other methods, as detailed in Tab.1. Our method outperformed others. As illustrated in Tab.2, we assessed user-perceived emotion levels when presented with videos labeled $(emo,level)$.
[…]
13.184 2.081 -11.000
12.250 2.911 -1.000
10.727 6.330 0.000
10.403 6.563 0.000
10.152 6.876 0.000
9.800 7.284 0.000
Table 3: Quantitative comparisons in ablation study.
Figure 11: Generation in ablation study.
4.4 Ablation Study.
We conducted ablation studies in multiple conditions to comprehensively demonstrate the contribution. Additionally, the absence of serial connection between two-level CDANs leads to poor performance. The incomplete methods shown in Fig. 11 lose the ability to express distinct emotional details, causing angry, surprised, and sad to resemble happy-like emotions, primarily by widening the mouth as the emotion level increases. Opt., OD and AS represents the mechanisms referred to Eq. 9 and Eq. 14 in LES Recon., Offset Decoder and Au Source respectively. We display progressively incorporating components in Tab. 3. More details are in the supplementary materials.
5 Conclusion
In this paper, we propose Linear Emotion Space (LES), a fine-grained emotion definition capturing facial unit actions and subtle details beyond AUs, along with a fine-grained emotion editing model, LES-Talker. We use 3DMM as an intermediate representation and propose Cross-Dimension Attention Net to guide the controllable deformation of the 3D model. We also design various key components to adapt multi-modal features. LES-Talker delivers high visual quality and fine-grained emotion editing, performing diverse tasks. Our theoretical definition of LES distinguishes the model and could serve as a foundation for future research.